### v4

@@ 1 Introduction
with a discount rate $r>0$, a running cost $f$, and an intervention cost $l$. Compared to regular control, impulse control is a more versatile mathematical framework as it incorporates discontinuities into the state processes. It also generalizes optimal stopping or the singular control problems, as impulse interventions can incur a nonzero fixed cost HT1983 . As a result, various problems in mathematical finance, engineering and management can be modeled as impulse control problems, including cash management constantinides1978existence , inventory controls Harrison1983 ; Sulem1986 , transaction cost in portfolio analysis Eastham1988 ; Morton1995 ; Korn1998 ; Korn1999 ; Bielecki2000 ; Oksendal2002 , insurance model Jeanblanc-Picque1995 ; Candenillas2006 , liquidity risk LyVath2007 , exchange rates Jeanblanc-Picque1993 ; Mundaca1998 ; Bertola2016 , real options Triantis1990 ; Mauer1994 ; BensoussanRoignant , energy market basei2019optimal , and so on.
@@
where $\mathcal{M}$ is a nonlocal operator such that $\mathcal{M}\psi(x)=\inf_{\xi\in\mathbb{R}}\{\psi(x+\xi)+l(\xi)\}$; see, for instance, bensoussan1984impulse ; oksendal2007applied for more details on the QVIs. Due to the very presence of this nonlocal operator and correspondingly the discontinuous nature of the controlled process, it takes delicate treatments of the value function, especially its regularity conditions, so that popular methods such as the smooth-fit principle can be applied; see for instance guo2009smooth ; davis2010impulse ; bayraktar2013impulse . One of the best known strategies would be the threshold-type $(d,D,U,u)$ derived for the one-dimensional cash management model in constantinides1978existence : the controller intervenes only if the state process exits the continuation region $(d,u)$, and in this case, the state is immediately adjusted to the closer threshold between $D$ and $U$; see also alvarez2004class for more explicitly solvable cases. It is particularly challenging to explicitly derive the value function and the corresponding optimal strategy under a generic impulse control framework.
@@
To numerically solve impulse control problems instead, several numerical methods for the corresponding HJB-QVIs have been developed. Many of these methods are based on iterative partial differential equation (PDE) frameworks. The most common route is through penalization combined with policy/Howard iteration: one solves a sequence of penalized HJB equations and iterates to a fixed point (see, e.g., bensoussan1984impulse ; reisinger2019penalty ; reisinger2020error ). Other widely used approaches include direct-control or semi-Lagrangian discretizations with policy iteration and matrix-monotonicity arguments azimzadeh2016weakly ; azimzadeh2018convergence , finite-difference (FD)/finite-element (FE) schemes that provide $L^{\infty}$-type error bounds for (non)coercive QVIs boulbrachene1998noncoercive ; boulbrachene2005pointwise , implicit time-stepping for finite-horizon formulations ieda2015implicit , and probabilistic representations via backward stochastic differential equations (BSDEs) that enable Monte-Carlo solvers for certain classes kharroubi2010backward . While effective for low dimensional state space, these methods face serious limitations in modern applications: grid-based PDE schemes suffer from the curse of dimensionality and delicate monotonicity/stencil requirements; penalty formulations introduce stiffness and parameter tuning (often degrading accuracy near free boundaries); FD/FE error analyses can hinge on regularity or coercivity not guaranteed in practice; and BSDE solvers may require nested simulation and restrictive jump structures.
@@
As an alternative way to bypass the aforementioned obstacles in traditional numerical methods, reinforcement learning (RL) has been widely studied in recent years to solve optimal stochastic control problems under the continuous setting. The essence of RL is to learn about the possibly unknown environment from either the historical trajectory data (offline) or the interactions (online), and then to solve a decision making problem according to some objective function. Therefore, it is crucial to balance between learning and decision-making, also known as the “exploration–exploitation” tradeoff. To encourage exploration, a randomized optimal control framework is necessary. For regular control problems, a common randomization scheme is through an entropy regularization over the policy; see, for instance, wang2020reinforcement . Due to the newly-introduced Shannon’s entropy term, the controller is encouraged to explore the entire trajectory space instead of only exploiting the optimal one under the deterministic setting. In wang2020reinforcement , a thorough analysis is conducted on a linear-quadratic problem, where the exploration–exploitation tradeoff is full characterized through the optimal randomized feedback strategy as a Gaussian distribution. The analysis of the entropy-regularized framework under a general setting is conducted in tang2022exploratory . Under this framework, model-free updates have been developed in continuous-time temporal-difference learning and $q$-learning jia2022policy ; jia2023q , where the control randomization provides a policy-gradient estimator and has a variance reduction effect in the meantime.
@@
To design learning algorithms for optimal stopping problems, a proper randomized framework essentially increases the frequency of receiving the otherwise extremely sparse reward signals associated with the stopping. One way to achieve that is to randomize the survival process $\{\mathbbm{1}\{\tau>t\}\}_{t\geq 0}$ for any given stopping time $\tau$. For instance, in dong2024randomized , the survival process is modeled as $\{e^{-\int_{0}^{t}\pi_{s}ds}\}_{t\geq 0}$ that is controlled by the nonnegative survival rate $\pi=\{\pi_{t}\}_{t\geq 0}$. With a running normalized entropy regularization $\mathcal{R}(\pi_{t})$, the randomized optimal stopping problem is solved via analyzing the associated semi-linear HJB equation. A corresponding offline RL algorithm is proposed. In dianetti2024exploratory , the survival process is modeled as $\{1-\xi_{t}\}_{t\geq 0}$ where $\xi=\{\xi_{t}\}_{t\geq 0}$ is a nondecreasing singular control process valued on $[0,1]$. With a cumulative relative entropy regularization on $\xi$, the randomized optimal stopping problem is formulated as a singular control problem whose state space is augmented with the current survival probability $y\in[0,1]$; its solution can be derived through its obstacle-type HJB equation. Correspondingly, two versions of RL algorithms, one model-based and the other model-free, are proposed to learn the decision boundary characterizing the optimal bang-bang control. In both dong2024randomized and dianetti2024exploratory , the respective randomized optimal stopping frameworks via the survival process are shown to converge to their classical counterparts as the parameter $\lambda$ modulating the randomization decreases to 0.
@@
Extending these randomization ideas to impulse control problems is nontrivial, as control policy concerns with not just one but a sequence of stopping times together with the associated jumps. In denkert2025control , a special case of impulse control problem is considered where the policy $\nu$ is modeled as a marked point process with an associated (uncontrolled) Poisson random measure; see kharroubi2010backward for more details. Instead of randomizing the survival processes for each stopping time, the randomization in denkert2025control is through a controlled intensity process that tilts the probability measure. Under the intensity policy, the policy gradient is derived and a corresponding actor-critic RL algorithm is then proposed.
@@
subject to $dY_{t}=b(Y_{t})dt+\sigma(Y_{t})dW_{t}$; see bensoussan1984impulse ; guo2009smooth . We propose a two-fold randomization scheme for impulse control problem. First, we randomize the nonlocal operator into $\mathcal{M}^{\lambda}$ for some $\lambda>0$ such that for any suitable test function $\phi$,
@@
where $\mathcal{R}(\pi_{t})$ is the normalized entropy as in dong2024randomized to encourage the exploration of intervention times. Given any $\bm{\lambda}=(\lambda_{1},\lambda_{2})\in(0,+\infty)^{2}$, the value function $\psi^{\bm{\lambda}}$ of the randomized impulse control problem is the fixed point for the compound operator $\mathcal{T}^{\bm{\lambda}}=T^{\lambda_{1}}\circ\mathcal{M}^{\lambda_{2}}$; the intensity-jump distribution pair $(\pi^{*},\mu^{*})$ denotes the corresponding optimal optimal policy if $\psi^{\bm{\lambda}}$ is attained at $(\pi^{*},\mu^{*})$.
@@
The iterative approach in the proof of existence for $\psi^{\bm{\lambda}}$ naturally gives rise to the corresponding RL algorithm, where the iterates correspond to the policy improvement steps whose convergence to the randomized value function is guaranteed by this existence proof. Under the same set of bounded assumptions, we can further derive a geometric convergence rate. We then design a temporal–difference (TD)-based algorithm. It provides an offline model-free learning procedure to solve the randomized impulse control problem. We implement and evaluate this RL algorithm on a classical linear model that has been widely explored in the impulse–control literature; see constantinides1978existence ; jeanblanc1993impulse ; guo2009smooth , for instance. The experiments empirically verify the convergence of the randomized problem to the classical one as $(\lambda_{1},\lambda_{2})$ approaches $(0,0)$. We also provide a sensitivity analysis with respect to the volatility parameter $\sigma$ in the linear model under the randomized setting. The exploration–exploitation trade off is demonstrated through the impact of $\sigma$ on value functions, intervention intensities as well as jump distributions.
@@ 2 Randomized Impulse Control Problem
Under Assumption 2.1, (2.1) admits a strong solution adapted to $\mathbb{F}$ such that $\mathbb{E}[X_{t}^{2}]<\infty$ for all $t\geq 0$; see, for instance, (karatzas1998brownian, , Theorem 2.9). In addition, we have the following technical lemma which is a direct result of applying Itô’s formula, Grönwall’s inequality and Cauchy-Schwarz inequality.
@@ 2.1 Review of the Classical Impulse Control Problem
In the literature of the classical impulse control problem without randomization (see, for instance, guo2009smooth ), the objective can be seen as the following optimization problem,
@@
Classical results bensoussan1984impulse ; guo2009smooth show that the value function $\psi$ in (2.3) is the unique viscosity solution of following Hamilton-Jacobi-Bellman (HJB) equation of the quasi-variational-inequality type,
@@
Apart from characterizing the optimal impulse control policy via (HJB-QVI), an equivalent characterization has also been analyzed in bensoussan1984impulse ; guo2009smooth via the following compound optimal stopping operator $\mathcal{T}$, where for any test function $\phi:\mathbb{R}\to\mathbb{R}$ that is bounded from below,
@@
subject to (2.1), where $\mathbb{T}$ denotes the set of $\mathbb{F}$-stopping times such that $\mathbb{P}(\tau<\infty)=1$ for any $\tau\in\mathbb{T}$, and $\mathbb{E}_{x}\left[\cdot\right]=\mathbb{E}\left[\cdot\biggl|X_{0-}=x\right]$. In particular, the value function $\psi$ in (2.3) is a fixed point for $\mathcal{T}$, that is, $\psi=\mathcal{T}\psi$ on $\mathbb{R}$. Correspondingly, from (2.8), we know that $\mathcal{T}\psi(x)$ defined by (classical) admits an optimal stopping time with respect to (2.1),
@@ 2.2 Randomization of Impulse Controls
Let us first re-examine the deterministic nature of the optimal classical impulse control policy in (2.8)–(2.9). and then proceed to the corresponding two-fold randomization scheme for impulse controls. The randomized operators introduced in Sections 2.2.1 and 2.2.2 are the key ingredients for the formulation of the randomized impulse control problem in Section 2.3.
@@ 2.2.1 Randomizing the Jump
For any $x\in\mathbb{R}$, let $\Phi_{x}$ denote the Gaussian distribution $\mathcal{N}(-x,1)$ with mean x-x and variance $1$. Intuitively, $\Phi_{x}$ denotes a simple randomized jump strategy at $x$ such that on average, the jump should send the state back to the origin.
@@ 2.2.2 Randomizing the Intervention Time
Similar to dong2024randomized , we borrow the terminology from Cox process jeanblanc2009mathematical with $\pi_{t}\geq 0$ being the intensity for any $t\geq 0$, and define the survival probability by time $t\geq 0$ as $p_{t}=\exp\left\{-\int_{0}^{t}\pi_{s}\mathrm{d}s\right\}$; or equivalently, the process $p=\{p_{t}\}_{t\geq 0}$ satisfies $p_{0}=1$ and

### v5

@@ 1 Introduction
with a discount rate $r>0$, a running cost $f$, and an intervention cost $l$. Compared to regular control, impulse control is a more versatile mathematical framework as it incorporates discontinuities into the state processes. It also generalizes optimal stopping or the singular control problems, as impulse interventions can incur a nonzero fixed cost Harrison and Taksar (1983). As a result, various problems in mathematical finance, engineering and management can be modeled as impulse control problems, including cash management Constantinides and Richard (1978), inventory controls Harrison et al. (1983); Sulem (1986), transaction cost in portfolio analysis Eastham and Hastings (1988); Morton and Pliska (1995); Korn (1998, 1999); Bielecki and Pliska (2000); Øksendal and Sulem (2002), insurance model Jeanblanc-Picqué and Shiryaev (1995); Cadenillas et al. (2006), liquidity risk Ly Vath et al. (2007), exchange rates Jeanblanc-Picqué (1993a); Mundaca and Oksendal (1998); Bertola et al. (2016), real options Triantis and Hodder (1990); Mauer and Triantis (1994); Bensoussan and Chevalier-Roignant (2019), energy market Basei (2019), and so on.
@@
where $\mathcal{M}$ is a nonlocal operator such that $\mathcal{M}\psi(x)=\inf_{\xi\in\mathbb{R}}\{\psi(x+\xi)+l(\xi)\}$; see, for instance, Bensoussan and Lions (1984); Øksendal and Sulem (2007) for more details on the QVIs. Due to the very presence of this nonlocal operator and correspondingly the discontinuous nature of the controlled process, it takes delicate treatments of the value function, especially its regularity conditions, so that popular methods such as the smooth-fit principle can be applied; see for instance Guo and Wu (2009); Davis et al. (2010); Bayraktar et al. (2013). One of the best known strategies would be the threshold-type $(d,D,U,u)$ derived for the one-dimensional cash management model in Constantinides and Richard (1978): the controller intervenes only if the state process exits the continuation region $(d,u)$, and in this case, the state is immediately adjusted to the closer threshold between $D$ and $U$; see also Alvarez (2004) for more explicitly solvable cases. It is particularly challenging to explicitly derive the value function and the corresponding optimal strategy under a generic impulse control framework.
@@
To numerically solve impulse control problems instead, several numerical methods for the corresponding HJB-QVIs have been developed. Many of these methods are based on iterative partial differential equation (PDE) frameworks. The most common route is through penalization combined with policy/Howard iteration: one solves a sequence of penalized HJB equations and iterates to a fixed point (see, e.g., Bensoussan and Lions (1984); Reisinger and Zhang (2019, 2020)). Other widely used approaches include direct-control or semi-Lagrangian discretizations with policy iteration and matrix-monotonicity arguments Azimzadeh and Forsyth (2016); Azimzadeh et al. (2018), finite-difference (FD)/finite-element (FE) schemes that provide $L^{\infty}$-type error bounds for (non)coercive QVIs Boulbrachene (1998, 2005), implicit time-stepping for finite-horizon formulations Ieda (2015), and probabilistic representations via backward stochastic differential equations (BSDEs) that enable Monte-Carlo solvers for certain classes Kharroubi et al. (2010). While effective for low dimensional state space, these methods face serious limitations in modern applications: grid-based PDE schemes suffer from the curse of dimensionality and delicate monotonicity/stencil requirements; penalty formulations introduce stiffness and parameter tuning (often degrading accuracy near free boundaries); FD/FE error analyses can hinge on regularity or coercivity not guaranteed in practice; and BSDE solvers may require nested simulation and restrictive jump structures.
@@
As an alternative way to bypass the aforementioned obstacles in traditional numerical methods, reinforcement learning (RL) has been widely studied in recent years to solve optimal stochastic control problems under the continuous setting. The essence of RL is to learn about the possibly unknown environment from either the historical trajectory data (offline) or the interactions (online), and then to solve a decision making problem according to some objective function. Therefore, it is crucial to balance between learning and decision-making, also known as the “exploration–exploitation” tradeoff. To encourage exploration, a randomized optimal control framework is necessary. For regular control problems, a common randomization scheme is through an entropy regularization over the policy; see, for instance, Wang et al. (2020). Due to the newly-introduced Shannon’s entropy term, the controller is encouraged to explore the entire trajectory space instead of only exploiting the optimal one under the deterministic setting. In Wang et al. (2020), a thorough analysis is conducted on a linear-quadratic problem, where the exploration–exploitation tradeoff is full characterized through the optimal randomized feedback strategy as a Gaussian distribution. The analysis of the entropy-regularized framework under a general setting is conducted in Tang et al. (2022). Under this framework, model-free updates have been developed in continuous-time temporal-difference learning and $q$-learning Jia and Zhou (2022, 2023), where the control randomization provides a policy-gradient estimator and has a variance reduction effect in the meantime.
@@
To design learning algorithms for optimal stopping problems, a proper randomized framework essentially increases the frequency of receiving the otherwise extremely sparse reward signals associated with the stopping. One way to achieve that is to randomize the survival process $\{\mathbbm{1}\{\tau>t\}\}_{t\geq 0}$ for any given stopping time $\tau$. For instance, in Dong (2024), the survival process is modeled as $\{e^{-\int_{0}^{t}\pi_{s}ds}\}_{t\geq 0}$ that is controlled by the nonnegative survival rate $\pi=\{\pi_{t}\}_{t\geq 0}$. With a running normalized entropy regularization $\mathcal{R}(\pi_{t})$, the randomized optimal stopping problem is solved via analyzing the associated semi-linear HJB equation. A corresponding offline RL algorithm is proposed. In Dianetti et al. (2024), the survival process is modeled as $\{1-\xi_{t}\}_{t\geq 0}$ where $\xi=\{\xi_{t}\}_{t\geq 0}$ is a nondecreasing singular control process valued on $[0,1]$. With a cumulative relative entropy regularization on $\xi$, the randomized optimal stopping problem is formulated as a singular control problem whose state space is augmented with the current survival probability $y\in[0,1]$; its solution can be derived through its obstacle-type HJB equation. Correspondingly, two versions of RL algorithms, one model-based and the other model-free, are proposed to learn the decision boundary characterizing the optimal bang-bang control. In both Dong (2024) and Dianetti et al. (2024), the respective randomized optimal stopping frameworks via the survival process are shown to converge to their classical counterparts as the parameter $\lambda$ modulating the randomization decreases to 0.
@@
Extending these randomization ideas to impulse control problems is nontrivial, as control policy concerns with not just one but a sequence of stopping times together with the associated jumps. In Denkert et al. (2025), a special case of impulse control problem is considered where the policy $\nu$ is modeled as a marked point process with an associated (uncontrolled) Poisson random measure; see Kharroubi et al. (2010) for more details. Instead of randomizing the survival processes for each stopping time, the randomization in Denkert et al. (2025) is through a controlled intensity process that tilts the probability measure. Under the intensity policy, the policy gradient is derived and a corresponding actor-critic RL algorithm is then proposed.
@@
subject to $dY_{t}=b(Y_{t})dt+\sigma(Y_{t})dW_{t}$; see Bensoussan and Lions (1984); Guo and Wu (2009). We propose a two-fold randomization scheme for impulse control problem. First, we randomize the nonlocal operator into $\mathcal{M}^{\lambda}$ for some $\lambda>0$ such that for any suitable test function $\phi$,
@@
where $\mathcal{R}(\pi_{t})$ is the normalized entropy as in Dong (2024) to encourage the exploration of intervention times. Given any $\bm{\lambda}=(\lambda_{1},\lambda_{2})\in(0,+\infty)^{2}$, the value function $\psi^{\bm{\lambda}}$ of the randomized impulse control problem is the fixed point for the compound operator $\mathcal{T}^{\bm{\lambda}}=T^{\lambda_{1}}\circ\mathcal{M}^{\lambda_{2}}$; the intensity-jump distribution pair $(\pi^{*},\mu^{*})$ denotes the corresponding optimal optimal policy if $\psi^{\bm{\lambda}}$ is attained at $(\pi^{*},\mu^{*})$.
@@
The iterative approach in the proof of existence for $\psi^{\bm{\lambda}}$ naturally gives rise to the corresponding RL algorithm, where the iterates correspond to the policy improvement steps whose convergence to the randomized value function is guaranteed by this existence proof. Under the same set of bounded assumptions, we can further derive a geometric convergence rate. We then design a temporal–difference (TD)-based algorithm. It provides an offline model-free learning procedure to solve the randomized impulse control problem. We implement and evaluate this RL algorithm on a classical linear model that has been widely explored in the impulse–control literature; see Constantinides and Richard (1978); Jeanblanc-Picqué (1993b); Guo and Wu (2009), for instance. The experiments empirically verify the convergence of the randomized problem to the classical one as $(\lambda_{1},\lambda_{2})$ approaches $(0,0)$. We also provide a sensitivity analysis with respect to the volatility parameter $\sigma$ in the linear model under the randomized setting. The exploration–exploitation trade off is demonstrated through the impact of $\sigma$ on value functions, intervention intensities as well as jump distributions.
@@ 2 Randomized Impulse Control Problem
Under Assumption 2.1, (2.1) admits a strong solution adapted to $\mathbb{F}$ such that $\mathbb{E}[X_{t}^{2}]<\infty$ for all $t\geq 0$; see, for instance, (Karatzas and Shreve, 1998, Theorem 2.9). In addition, we have the following technical lemma which is a direct result of applying Itô’s formula, Grönwall’s inequality and Cauchy-Schwarz inequality.
@@ 2.1 Review of the Classical Impulse Control Problem
In the literature of the classical impulse control problem without randomization (see, for instance, Guo and Wu (2009)), the objective can be seen as the following optimization problem,
@@
Classical results Bensoussan and Lions (1984); Guo and Wu (2009) show that the value function $\psi$ in (2.3) is the unique viscosity solution of following Hamilton-Jacobi-Bellman (HJB) equation of the quasi-variational-inequality type,
@@
Apart from characterizing the optimal impulse control policy via (HJB-QVI), an equivalent characterization has also been analyzed in Bensoussan and Lions (1984); Guo and Wu (2009) via the following compound optimal stopping operator $\mathcal{T}$, where for any test function $\phi:\mathbb{R}\to\mathbb{R}$ that is bounded from below,
@@
subject to (2.1), where $\mathbb{T}$ denotes the set of $\mathbb{F}$-stopping times such that $\mathbb{P}(\tau<\infty)=1$ for any $\tau\in\mathbb{T}$, and $\mathbb{E}_{x}\left[\cdot\right]=\mathbb{E}\left[\cdot|X_{0-}=x\right]$. In particular, the value function $\psi$ in (2.3), and $\mathcal{T}\psi$ as defined in (classical), both solve the same (HJB-QVI) with $\mathcal{M}\phi$ replaced by $\mathcal{M}\psi$. It can be shown that $\psi$ is a fixed point for $\mathcal{T}$, that is, $\psi=\mathcal{T}\psi$ on $\mathbb{R}$. Correspondingly, from (2.8), we know that $\mathcal{T}\psi(x)$ defined by (classical) admits an optimal stopping time with respect to (2.1),
@@ 2.2 Randomization of Impulse Controls
The deterministic nature of the optimal classical impulse control policy in (2.8)–(2.9) lies at that both the intervention time and the jump are deterministic functions of the state process. This deterministic “wait–then–intervene” structure is natural when the environment $(b,\sigma,f,l)$ is fully specified, but it becomes restrictive when some or all of these components are only partially known and must be inferred from data. In such settings, classical grid-based HJB–QVI solvers are difficult to deploy, whereas RL methods are more natural.
Inspired by (classical), we propose a two-fold randomization scheme for impulse controls. The classical policy induces sparse and highly structured feedback: interventions occur only at a finite set of deterministic boundary points and the chosen jumps bring the state process to corresponding optimal positions, which yields limited exploration and weak learning signals. By contrast, allowing randomness in both the jump distribution and the intervention time produces smoother, more frequent feedback and an intrinsic exploration mechanism; this is crucial for balancing exploration and exploitation in the subsequent RL framework.
The following two randomized operators are the key ingredients for the formulation of the randomized impulse control problem in Section 2.3.
@@ 2.2.1 Randomizing the Jump
Our Assumption 2.3 on the structure of the cost functions ensures that the classical value function is approximately “V-shaped”, where the origin is near the unique minimum and inside the continuation region. Inspired by this observation, we adopt a reference jump distribution so that the state process is driven towards the origin where the value function is near its minimum.
In specific, for any $x\in\mathbb{R}$, let $\Phi_{x}$ denote the Gaussian distribution $\mathcal{N}(-x,1)$ with mean x-x and variance $1$. Intuitively, $\Phi_{x}$ denotes a simple randomized jump strategy at $x$ such that on average, the jump should send the state back to the origin.
@@ 2.2.2 Randomizing the Intervention Time
Similar to Dong (2024), we borrow the terminology from Cox process Jeanblanc et al. (2009) with $\pi_{t}\geq 0$ being the intensity for any $t\geq 0$, and define the survival probability by time $t\geq 0$ as $p_{t}=\exp\left\{-\int_{0}^{t}\pi_{s}\mathrm{d}s\right\}$; or equivalently, the process $p=\{p_{t}\}_{t\geq 0}$ satisfies $p_{0}=1$ and